OpenStudy (anonymous):

how can this be simplfied? 9 +/- sq. root of 61/2

OpenStudy (saifoo.khan):

\[\Large 9\pm \sqrt{\frac{61}{2}}\]

OpenStudy (anonymous):

\[\frac{9\pm\sqrt{61}}{2}\]?

OpenStudy (anonymous):

if so , the answer is no

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

61 is prime, so there is nothing you can do with \[\sqrt{61}\]

OpenStudy (anonymous):

ok, thanks

OpenStudy (anonymous):

yw